OpenStudy (anonymous):

indefinite integral cos(t^1/2)/(t^1/2)

OpenStudy (anonymous):

\[\int\limits_{?}^{?} \cos \sqrt{t} / ( \sqrt{t})\]

OpenStudy (xapproachesinfinity):

did you consider U substitution for this

OpenStudy (xapproachesinfinity):

it will still involve the integral by part

OpenStudy (xapproachesinfinity):

I guess

OpenStudy (xapproachesinfinity):

@javikun1

OpenStudy (xapproachesinfinity):

let U=root(t) then du=-1/t^2dt, dt=-t^2du

OpenStudy (xapproachesinfinity):

Any progress?

OpenStudy (xapproachesinfinity):

@Zarkon, what do you suggest?

OpenStudy (anonymous):

i dont understand what you did

OpenStudy (xapproachesinfinity):

I did U substitution. Do you know of it? Actually I thought back about it. I did a mistake there du= (1/2root(x))dx not what I did above.